Read and analyze the following situations and supply the values of the following variables (if

there is any). On the third column, write known on the space provided if the situation gives

or can compute the value of the variance, otherwise write unknown. Identify also the formula

to be used to estimate the standard error of the mean by writing the symbols or when the

population variance is known and sx when the population variance is unknown.


Situation:

Given the population mean of 12, and a sample

standard deviation of 3 in a sample size of 125


Given:

n=

m=

s=

Answer:

Standard Error Formula:


1
Expert's answer
2022-03-21T13:51:13-0400

n=125

m=12

s=3

"SE=\frac{\sigma}{\sqrt{n}}=\frac{3}{\sqrt{125}}=0.27"